After testing for nitrates, the three professors decide to dive. Deep in the ocean, they pull out their favorite thermometer find that the water is -7°C What is the concentration of salt (NaCI, 58.4 g/mol) in molarity that is preventing the water from freezing? (density of seawater = 1.03 g/mL and K, 1.86°C/m)?
Answer #1

let m be the molality
and M be the molarity
use,
(delta Tf) = i*Kf*m
i for NaCl is 2
so,
0 - (-0.7) = 2*1.86*m
0.7 = 3.72*m
m = 0.188 m

Mb = molar mass of NaCl
= 58.4 g/mol
use,
m = (1000*M)/(M*Mb - 1000*d)
0.188 = (1000*M)/(58.4*M - 1000*1.03)
0.188 = (1000*M)/(58.4*M -1030)
11.0*M - 193.6 = 1000*M
989*M = 193.6
M = 0.196 M

Answer : 0.196 M